OpenStudy (miss.rose):

What is 7/10 - 5/8? Help and explain! Please show on a number line! I will give medal and fan! @zoeeh127 @rebeccaxhawaii

OpenStudy (anonymous):

3/40. Here is a link to a really simple explanation (a lot better than I could do). https://www.mathsisfun.com/fractions_subtraction.html

OpenStudy (miss.rose):

Thank you @Allenismyname That was a big help. But does it show how to do it on a number line?

OpenStudy (anonymous):

Not on a number line. It is easier to just perform the equation itself, plus I am not solid on the logistics of subtracting different fractions on a number line

OpenStudy (miss.rose):

I agree. But my assignments wants me to do it on a number line. Sadly.

rebeccaxhawaii (rebeccaxhawaii):

what do you need help with ?

OpenStudy (miss.rose):

7/10 - 5/8 = ? And can you help me put it on a number line. I know I am asking for a lot but I just really need help.

rebeccaxhawaii (rebeccaxhawaii):

\[\frac{ 7 }{ 10 }=\frac{ 28 }{ 40} \]

rebeccaxhawaii (rebeccaxhawaii):

\[\frac{ 5 }{ 8 }=\frac{ 25}{ 40 }\]

rebeccaxhawaii (rebeccaxhawaii):

so 28/40-25/40= what ?

OpenStudy (frozen15):

7/10 - 5/8 = 0.075 = 75/1000 = 3/40 sorry but I don't know how to make a number line on OS , but this is the answer ^

OpenStudy (miss.rose):

@Frozen15 just click the button that says, "Draw" it is the middle one at the bottom of the comment section.

OpenStudy (frozen15):

oh ok thanks!

rebeccaxhawaii (rebeccaxhawaii):



rebeccaxhawaii (rebeccaxhawaii):

is that all your questions ?

OpenStudy (miss.rose):

Um... Yes. Thank you both!!!! I wish I could give you both medals but I can only do one so I am going to medal @rebeccaxhawaii Nothing personal @Frozen15 it's just she helped me more. But I will fan you!

rebeccaxhawaii (rebeccaxhawaii):

ok thanks

OpenStudy (frozen15):

I have a idea! @rebeccaxhawaii can medal me and u can medal her?

OpenStudy (frozen15):

just suggestion :) thanks for fanning me

OpenStudy (miss.rose):

Your welcome! I might need you help later today so I will just tag you. Is that okay?

OpenStudy (frozen15):

that's fine! if I'm on , I will help :)

OpenStudy (miss.rose):

Okay. Thank you, again. Well, I am going to close this question. Bye!